OpenStudy (anonymous):

Estimate the difference 4 19/20 - 1 1/8 A. 3 B. 4 C. 3 1/2 or D. 3 1/4

OpenStudy (anonymous):

Use the same problem solver as the last question. 4+1=?

OpenStudy (anonymous):

oops -*

OpenStudy (anonymous):

5

OpenStudy (anonymous):

no i meant minus

OpenStudy (anonymous):

4-1=?

OpenStudy (anonymous):

The answer is B.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

wow. u just went against CoC -.-

OpenStudy (anonymous):

@ElloLovee follow CoC next time plz

OpenStudy (anonymous):

She was right tho. cuz 4-1=3 19/20 and 1/8 round up to 1 1+3=4

OpenStudy (anonymous):

Why would you tell me? you have to work me though the problem not just tell me!

OpenStudy (anonymous):

So the answer is B

OpenStudy (anonymous):

yes good job